the sum of the ages of a father and a son is 56. Four years ago the father was 3 times as old as the son. find the age of each
Answer by addingup(3677)

f+s= 56 if we subtract s from both sides:
f= 56-s We'll use this value for f in the next equation
----------------
f-4= 3s Substitute for f:
(56-s)-4= 3s
56-s-4= 3s add s on both sides and do subtraction on left:
52= 4s Divide both sides by 4
s= 13 The son is 13 and the father is:
56-13= 43
J